-36 + 2w = -8w + w

4p - 10 = p + 3p - 2p

Answer by jim_thompson5910(35256)

-36+2w=-8w+w



-36+2w=-7w



2w=-7w+36



2w+7w=36



9w=36



w=(36)/(9)



w=4

So the solution is w = 4

-------------------------------------------------------

4p-10=p+3p-2p



4p-10=2p



4p=2p+10



4p-2p=10



2p=10



p=(10)/(2)



p=5

So the solution is p = 5
-------------------------------------------------------------------------------------------------
If you need more help, email me at jim_thompson5910@hotmail.com

Also, please consider visiting my website: http://www.freewebs.com/jimthompson5910/home.html and making a donation. Thank you

Jim

-------------------------------------------------------------------------------------------------